# Graph Knowledge Distillation to Mixture of Experts

### v1

@@ Abstract
In terms of accuracy, Graph Neural Networks (GNNs) are the best architectural choice for the node classification task. Their drawback in real-world deployment is the latency that emerges from the neighbourhood processing operation. One solution to the latency issue is to perform knowledge distillation from a trained GNN to a Multi-Layer Perceptron (MLP), where the MLP processes only the features of the node being classified (and possibly some pre-computed structural information). However, the performance of such MLPs in both transductive and inductive settings remains inconsistent for existing knowledge distillation techniques. We propose to address the performance concerns by using a specially-designed student model instead of an MLP. Our model, named Routing-by-Memory (RbM), is a form of Mixture-of-Experts (MoE), with a design that enforces expert specialization. By encouraging each expert to specialize on a certain region on the hidden representation space, we demonstrate experimentally that it is possible to derive considerably more consistent performance across multiple datasets.
@@ 1 Introduction
Graphs can be used to encode the dependencies between data samples. The impressive performance of Graph Neural Networks (GNNs) shows that taking into account the structural information increases the quality of prediction on tasks like product prediction on co-purchasing graphs or paper category prediction on citation graphs (Kipf & Welling, 2016; Hamilton et al., 2017). However, despite the potential accuracy improvements of GNNs, multi-layer perceptrons (MLPs) remain preferable to graph neural networks for many large-scale industrial applications. This is due to the fundamental inefficiency of GNNs, with scalability limitations making deployment challenging (Zhang et al., 2020a; Jia et al., 2020; Zheng et al., 2022). GNNs operate in layers, and each layer requires the processing of a neighbourhood of the node in order to compute the prediction. For example, evaluating the prediction for a single node with an $L$-layer GNN requires processing at least every node in the $L$-hop neighbourhood. For real-world graphs, involving millions of nodes, the $L$-hop neighbourhood can be very large, leading to resource intensive operations (Jin et al., 2021). Even if we only sample a subset of the neighbours at each layer, the receptive field for the node can grow very rapidly. By contrast, forming a prediction for a single node with an $L$-layer MLP requires processing only the features of that node. This can be scaled and deployed efficiently, and parallelization is straightforward.
@@
  • We provide experimental evidence that simply increasing the size of a student MLP does not consistently lead to better performance.
  • To address this issue, we propose a Routing-by-Memory (RbM) model with enforced expert specialization.
  • We show that our approach outperforms baselines on datasets of different sizes, and we demonstrate that the proposed approach outperforms an ensemble model and a standard MoE.
@@ 2.2 Mixture-of-Experts
A Sparse Mixture-of-Experts (MoE) model is a weighted combination of similarly structured models with dynamically computed weights (Shazeer et al., 2016; Gross et al., 2017; Zhang et al., 2021a; Li et al., 2022; Dryden & Hoefler, 2022; Chi et al., 2022; Komatsuzaki et al., 2022; Pavlitska et al., 2023). For any sample, only a small portion of the experts have non-zero weights. This allows us to increase the model learning capacity without significantly inflating the processing costs (Fedus et al., 2022). An MoE can also be used as a layer inside a larger model (Qu et al., 2022; Yan & Li, 2023). In most implementations, the computation of the expert weights, referred to as routing, is perforrmed by a separate neural network (a policy network) of a size comparable to an expert. Based on the MoE input, it produces weights for the experts. Because both the policy network and expert networks are trained simultaneously, there is a danger of under-utilization, and consequently under-training, of some experts (Krishnamurthy et al., 2023). This is commonly mitigated by injecting noise and designing loss functions that even out the utilization.
An alternative routing scheme, designed to prevent routing inconsistencies, involves pairing each expert network with an embedding vector. Instead of predicting the expert weights directly, the policy network then aims to project the input sample into the embedding space. The weights are derived from the distances to the expert embeddings (Gross et al., 2017; Zhang et al., 2021a; Chi et al., 2022; Li et al., 2022; Yan & Li, 2023; Qu et al., 2022). Initially, the Euclidean distance was used to route to the single closest expert (Gross et al., 2017; Zhang et al., 2021a). However, this causes computational stability issues when multiple experts are used (Qu et al., 2022). Using dot-product similarity (Lample et al., 2019; Fedus et al., 2022) instead of Euclidean distance typically leads to representation collapse. Chi et al. (2022) used cosine similarity to avoid the collapse, and this is currently used in many implementations (Li et al., 2022; Yan & Li, 2023). Zhang et al. (2021a) propose moving the expert embeddings into the input space in order to encourage expert specialization. This leads to each expert specialising on the area around an embedding vector.
@@ 3 Background
We denote a graph by $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$, where $\mathcal{V}$ is a set of $N$ nodes, $\mathcal{E}$ is a set of edges between nodes, and $X\in\operatorname{\mathbb{R}}^{N\times d}$ represents a matrix with each row being a vector of $d$ node features associated with the corresponding node. For the node classification task, with $C$ classes, we use a label matrix, $Y\in\{0,1\}^{N\times C}$, with each row containing a one-hot encoded class label. The superscript $L$ denotes the labelled nodes of the graph and the superscript $U$ denotes unlabelled nodes, i.e., $\mathcal{V}^{L}$, $X^{L}$, $Y^{L}$ are, respectively, the labeled nodes, their node features, and the one-hot class labels.
@@ 4 Methodology
Figure 1: (a) An overview of the overall training framework. A teacher GNN is trained on the graph and provides targets for Knowledge Distillation (KD) (9) and Knowledge-Aware Reliable Distillation (KRD) (10) losses. A Mixture-of-Experts student is trained on the node features and positional encoding (see Section 4.2). (b) We use three additional losses to adjust the internal representations of the model, as the embeddings we use for routing (see Section 4.1). We provide schematic representations of these losses to aid intuitive understanding. Commitment loss (6) pulls representations closer to embeddings (highlighted in blue). Self-similarity loss (7) prevents collapse of representations. Load balance loss (8) helps to move borderline representations towards embedding of the less populated experts.
@@ 4.1 Spatial routing by memory
$h_{l}=\sum_{i=1}^{E}G(h_{l-1})_{i}f_{i}(h_{l-1}),$ (1)
where $G:\operatorname{\mathbb{R}}^{d^{\prime}}\rightarrow[0,1]^{E}$ is a policy network that produces routing coefficients, $f_{i}:\operatorname{\mathbb{R}}^{d^{\prime}}\rightarrow\operatorname{\mathbb{R}}^{d^{\prime\prime}}$ is an expert, $E$ is the number of experts, and $h_{l-1}\in\operatorname{\mathbb{R}}^{d^{\prime}}$ is the input hidden representation of the datapoint, $x\in\operatorname{\mathbb{R}}^{d}$, emerging from the $(l{-}1)$-th layer. For the first layer, the hidden representation is the input, i.e., $h_{0}=x$.
Li et al. (2022) introduce a routing scheme that uses a set of embeddings, $Q_{MoE}\in\operatorname{\mathbb{R}}^{E\times{d_{e}}}$, with each embedding being associated with a particular expert. The weight assigned to each expert for a given input at layer $l$ is determined by measuring the cosine distance between the projection of the input vector $h_{l-1}$ and the expert’s embedding. The policy network routes the input vector to the $k$ nearest experts:
$G_{MoE}(h)=\operatorname*{softmax}\left(\operatorname{Top}_{k}\left(\frac{Q^{MoE}Wh}{\|Q^{MoE}\|\|Wh\|}\right)\right)\,.$ (2)
Here $W\in\operatorname{\mathbb{R}}^{d_{e}\times{d^{\prime}}}$ is a trainable projection matrix, and the operation $\operatorname{Top}_{k}(\cdot)$ is a one-hot embedding that sets all elements in the output vector to zero except for the elements with the largest $k$ values. Chi et al. (2022) demonstrate that using a policy network of this form results in a more even distribution of token projections over the projection space, leading to improved performance.
While Li et al. (2022) provide analysis demonstrating that $G_{MoE}(\cdot)$ provides a degree of specialization for expert networks, we found it insufficient (see the experimental results in Section 5.5). We therefore enforce experts’ local specialization by setting the expert embeddings $Q^{RbM}\in\operatorname{\mathbb{R}}^{E\times{d^{\prime}}}$ to vectors in the input space, rather than projecting to a separate space. We achieve this by setting:
$G_{RbM}(h)=\operatorname*{softmax}\left(\operatorname{Top}_{k}\left(\frac{\operatorname{sg}[Q^{RbM}]h}{\|\operatorname{sg}[Q^{RbM}]\|\|h\|}\right)\right),$ (3)
v where $\operatorname{sg}[\cdot]$ is a stop gradient function. In our approach, each expert embedding vector is positioned at the center of an area in the representation space that the expert is specialising on. We use cosine similarity and interpolation between multiple experts.
@@
$Q^{RbM}_{i}(t+1)=\lambda Q^{RbM}_{i}(t)+(1-\lambda)\!\!\!\!\!\!\sum_{\begin{subarray}{c}j=1\\ G_{RbM}(h_{j})_{i}\neq 0\end{subarray}}^{B}\!\!\!\!\!\!\operatorname*{softmax}_{j}(G_{RbM}(h_{j})_{i})\,h_{j}\,,$ (4)
where $\lambda$ is a momentum coefficient hyperparameter. We compute this sum only for the vectors that are routed to the expert. Applying the softmax over the batch dimension improves performance.
@@
where $s\in\operatorname{\mathbb{R}}$ is a learnable output scaler, $att_{i}\in\operatorname{\mathbb{R}}^{d}$ is a learnable input attention vector, and $\scalerel\mathbin{\scalerel*{\bm{\odot}}{\circ}}$ denotes element-wise multiplication. Note that the attention is applied only to the expert input and not to the router input. This routing approach is inspired by the technique presented by Zhang et al. (2021a).
@@
$Loss_{VQ}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{E}G_{RbM}(h_{i})_{j}\frac{\operatorname{sg}[Q^{RbM}_{j}]h_{i}}{\|\operatorname{sg}[Q^{RbM}_{j}]\|\|h_{i}\|}\,,$ (6)
@@
$Loss_{SS}=\frac{1}{E^{2}}\sum_{i=1}^{E}\sum_{j=1}^{E}\frac{\operatorname{sg}[Q^{RbM}_{j}]Q^{RbM}_{i}}{\|\operatorname{sg}[Q^{RbM}_{j}]\|\|Q^{RbM}_{i}\|}\,.$ (7)
@@
$Loss_{LB}=\frac{Var(Load)}{Mean(Load)^{2}},\quad\quad\textrm{for}\quad Load=\sum_{i=1}^{B}G_{RbM}(h_{i}).$ (8)
Here $Load\in\operatorname{\mathbb{R}}^{E}$ is the vector of unnormalized expert utilization. While expert specialization does not require load balancing, we adopt this loss to redistribute input vectors that lie in almost equal proximity to multiple experts. This loss encourages representations to move closer to the least utilised expert embeddings.
@@ 4.2 Knowledge Distillation
$Loss_{KD}=\frac{\lambda}{\left|\mathcal{V}^{L}\right|}\sum_{v\in\mathcal{V}^{L}}CE(\hat{y}_{v},y_{v})+\frac{1-\lambda}{\left|\mathcal{V}\right|}\sum_{v\in\mathcal{V}}KL(\hat{y}_{v},\hat{y}^{\prime}_{v})\,.$ (9)
Here $\lambda$ is a hyperparameter that controls how much the knowledge distillation training relies on the soft labels from the teacher GNN, $CE(\cdot,\cdot)$ denotes cross-entropy, and $KL(\cdot,\cdot)$ is a KL-divergence (Zhang et al., 2021b). While the supervised part of the loss uses only the labeled part of the graph, the KL-divergence is computed over the entire set of nodes.
@@
In order to additionally leverage the graph structure, we employ the knowledge-based sampling technique of KRD (Wu et al., 2023). Given a target node $v$, we sample nodes from its neighbourhood, $\mathcal{N}(v)$, updating the sampling distribution after every training step based on how well the student model is predicting each node. and include a training loss that strives to match the class distribution predicted by the student with the teacher-provided soft labels of each sampled node:
$Loss_{KRD}=\frac{1-\lambda}{\left|\mathcal{V}\right|}\sum_{v\in\mathcal{V}}\operatorname{\mathbb{E}}_{u\sim\mathcal{N}(v)}KL(\hat{y}_{v},\hat{y}^{\prime}_{u}).$ (10)
The final objective is a combination of the cross-entropy loss, the distillation loss, the KRD loss, and embedding losses for every MoE layer of the model. With a model with $L$ MoE layers, we introduce weights $\alpha$, $\beta$, and $\gamma$ to determine the influences of the associated embedding losses:

### v2

@@ Abstract
In terms of accuracy, Graph Neural Networks (GNNs) are the best architectural choice for the node classification task. Their drawback in real-world deployment is the latency that emerges from the neighbourhood processing operation. One solution to the latency issue is to perform knowledge distillation from a trained GNN to a Multi-Layer Perceptron (MLP), where the MLP processes only the features of the node being classified (and possibly some pre-computed structural information). However, the performance of such MLPs, in both transductive and inductive settings, remains inconsistent for existing knowledge distillation techniques. We propose to address the performance concerns by using a specially-designed student model instead of an MLP. Our model, named Routing-by-Memory (RbM), is a form of Mixture-of-Experts (MoE), with a design that enforces expert specialization. By encouraging each expert to specialize on a certain region on the hidden representation space, we demonstrate experimentally that it is possible to derive considerably more consistent performance across multiple datasets. Code available at https://github.com/Rufaim/routing-by-memory.
@@ 1 Introduction
Graphs can be used to encode the dependencies between data samples. The impressive performance of Graph Neural Networks (GNNs) shows that taking into account the structural information increases the quality of prediction on tasks like product prediction on co-purchasing graphs or paper category prediction on citation graphs (Kipf & Welling, 2016; Hamilton et al., 2017). However, despite the potential accuracy improvements of GNNs, multi-layer perceptrons (MLPs) remain preferable to graph neural networks for many large-scale industrial applications. This is due to the fundamental inefficiency of GNNs, with scalability limitations making deployment challenging (Zhang et al., 2020a; Jia et al., 2020; Zheng et al., 2022). GNNs operate in layers, and each layer requires the processing of a neighbourhood of the node in order to compute the prediction. For example, evaluating the prediction for a single node with an $L$-layer GNN requires processing at least every node in the $L$-hop neighbourhood. For real-world graphs, involving millions of nodes, the $L$-hop neighbourhood can be very large, leading to resource intensive operations (Jin et al., 2021). Even if we only sample a subset of the neighbours at each layer, the receptive field for the node can grow very rapidly. This can lead to a latency performance bottleneck in high-load systems. This is especially the case in situations where the graph is distributed over multiple servers. Fetching the entire $L$-hop neighbourhood can result in greatly increased latency (Zheng et al., 2022). By contrast, forming a prediction for a single node with an $L$-layer MLP requires processing only the features of that node. This naturally eliminates the latency problems associated with additional node fetching. Therefore it can be scaled and deployed efficiently, and parallelization is straightforward.
@@
  • Ours is the first work to propose the use of a student Mixture-of-Experts (MoE) model for the distillation of a GNN.
  • We propose a Routing-by-Memory (RbM) model that differs from, and outperforms, a standard sparse MoE. We introduce important adaptations to a routing system previously proposed by Zhang et al. (2021a) for routing to a single expert in a computer vision setting. We allow for routing to multiple experts and use a different distance.
  • During training of the proposed MoE, we introduce several loss terms to encourage better clustering of representations and improved expert specialization.
  • We conduct multiple novel experiments, demonstrating that the proposed approach consistently outperforms both (i) enlarged MLP students; and (ii) ensembles or sparse MoEs.
@@ 2.2 Mixture-of-Experts
A Sparse Mixture-of-Experts (MoE) model is a weighted combination of similarly structured models with dynamically computed weights (Shazeer et al., 2016; Gross et al., 2017; Zhang et al., 2021a; Li et al., 2022; Dryden & Hoefler, 2022; Chi et al., 2022; Komatsuzaki et al., 2022; Pavlitska et al., 2023). For any sample, only a small portion of the experts have non-zero weights. This allows us to increase the model learning capacity without significantly inflating the processing costs (Fedus et al., 2022). An MoE can also be used as a layer inside a larger model (Qu et al., 2022; Yan & Li, 2023). In most implementations, the computation of the expert weights, referred to as routing, is performed by a separate neural network (a policy network) of a size comparable to an expert. Based on the MoE input, it produces weights for the experts. Because both the policy network and expert networks are trained simultaneously, there is a danger of under-utilization, and consequently under-training, of some experts (Krishnamurthy et al., 2023). This is commonly mitigated by injecting noise and designing loss functions that even out the utilization.
An alternative routing scheme, designed to prevent routing inconsistencies, involves pairing each expert network with an embedding vector. Instead of predicting the expert weights directly, the policy network then aims to project the input sample into the embedding space. The weights are derived from the distances to the expert embeddings (Gross et al., 2017; Zhang et al., 2021a; Chi et al., 2022; Li et al., 2022; Yan & Li, 2023; Qu et al., 2022). Initially, the Euclidean distance was used to route to the single closest expert (Gross et al., 2017; Zhang et al., 2021a). However, this causes computational stability issues when multiple experts are used (Qu et al., 2022). Using dot-product similarity (Lample et al., 2019; Fedus et al., 2022) instead of Euclidean distance typically leads to representation collapse. Chi et al. (2022) used cosine similarity to avoid the collapse, and this is currently used in many implementations (Li et al., 2022; Yan & Li, 2023). Zhang et al. (2021a) propose moving the expert embeddings into the layer’s input space in order to encourage expert specialization. In their scheme, the input vector is always routed to a single expert, the one with the closest embedding vector. Each expert embedding is updated by calculating a moving average over the input embedding vectors that are routed to that expert. This leads to each expert specialising on the area of the input space around its embedding.
@@ 3 Background
We denote a graph by $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}$ is a set of $N$ nodes, $\mathcal{E}$ is a set of edges between nodes, and $\mathbf{X}\in\operatorname{\mathbb{R}}^{N\times d}$ represents a matrix with each row being a vector of $d$ node features associated with the corresponding node. For the node classification task, with $C$ classes, we use a label matrix, $\mathbf{Y}\in\{0,1\}^{N\times C}$, with each row containing a one-hot encoded class label. The superscript $L$ denotes the labelled nodes of the graph and the superscript $U$ denotes unlabelled nodes, i.e., $\mathcal{V}^{L}$, $\mathbf{X}^{L}$, $\mathbf{Y}^{L}$ are, respectively, the labeled nodes, their node features, and the one-hot class labels.
@@ 4 Methodology
Figure 1: (a) An overview of the overall training framework. A teacher GNN is trained on the graph and provides targets for Knowledge Distillation (KD) (9) and Knowledge-Aware Reliable Distillation (KRD) (12) losses. A Mixture-of-Experts student is trained on the node features and positional encoding (see Section 4.2). (b) We use three additional losses to adjust the internal representations of the model, as the embeddings we use for routing (see Section 4.1). We provide schematic representations of these losses to aid intuitive understanding. Commitment loss (6) pulls representations closer to embeddings (highlighted in blue). Self-similarity loss (7) prevents collapse of representations. Load balance loss (8) helps to move borderline representations towards the embeddings of the less populated experts.
@@ 4.1 Spatial routing by memory
$\mathbf{h}_{l}=\sum_{i=1}^{E}G(\mathbf{h}_{l-1})_{i}f_{i}(\mathbf{h}_{l-1}),$ (1)
where $G:\operatorname{\mathbb{R}}^{d^{\prime}}\rightarrow[0,1]^{E}$ is a policy network that produces routing coefficients, $f_{i}:\operatorname{\mathbb{R}}^{d^{\prime}}\rightarrow\operatorname{\mathbb{R}}^{d^{\prime\prime}}$ is an expert, $E$ is the number of experts, and $\mathbf{h}_{l-1}\in\operatorname{\mathbb{R}}^{d^{\prime}}$ is the input hidden representation of the datapoint, $\mathbf{x}\in\operatorname{\mathbb{R}}^{d}$, emerging from the $(l{-}1)$-th layer. For the first layer, the hidden representation is the input, i.e., $\mathbf{h}_{0}=\mathbf{x}$.
Li et al. (2022) introduce a routing scheme that uses a set of embeddings, $\mathbf{Q}^{MoE}\in\operatorname{\mathbb{R}}^{E\times{d_{e}}}$, with each embedding being associated with a particular expert. The weight assigned to each expert for a given input at layer $l$ is determined by measuring the cosine distance between the projection of the input vector $h_{l-1}$ and the expert’s embedding. The policy network routes the input vector to the $k$ nearest experts:
$G_{MoE}(\mathbf{h})=\operatorname*{softmax}\left(\operatorname{Top}_{k}\left(\frac{\mathbf{Q}^{MoE}\mathbf{W}\mathbf{h}}{\|\mathbf{Q}^{MoE}\|\|\mathbf{W}\mathbf{h}\|}\right)\right)\,.$ (2)
Here $\mathbf{W}\in\operatorname{\mathbb{R}}^{d_{e}\times{d^{\prime}}}$ is a trainable projection matrix, and the operation $\operatorname{Top}_{k}(\cdot)$ is a one-hot embedding that sets all elements in the output vector to zero except for the elements with the largest $k$ values. Chi et al. (2022) demonstrate that using a policy network of this form results in a more even distribution of token projections over the projection space, leading to improved performance.
While Li et al. (2022) provide analysis demonstrating that $G_{MoE}(\cdot)$ provides a degree of specialization for expert networks, we found it insufficient (see the experimental results in Section 5.5). We therefore enforce experts’ local specialization by setting the expert embeddings $\mathbf{Q}^{RbM}\in\operatorname{\mathbb{R}}^{E\times{d^{\prime}}}$ to vectors in the input space, rather than projecting to a separate space. We achieve this by setting:
$G_{RbM}(h)=\operatorname*{softmax}\left(\operatorname{Top}_{k}\left(\frac{\operatorname{sg}[\mathbf{Q}^{RbM}]\mathbf{h}}{\|\operatorname{sg}[\mathbf{Q}^{RbM}]\|\|\mathbf{h}\|}\right)\right),$ (3)
where $\operatorname{sg}[\cdot]$ is a stop gradient function. In our approach, each expert embedding vector is positioned at the center of an area in the representation space that the expert is specialising on. We use cosine similarity and interpolation between multiple experts (see Figure 3).
@@
$\mathbf{Q}^{RbM}_{i}(t+1)=\hat{\lambda}(t)\mathbf{Q}^{RbM}_{i}(t)+(1-\hat{\lambda}(t))\!\!\!\!\!\!\sum_{\begin{subarray}{c}j=1\\ G_{RbM}(\mathbf{h}_{j})_{i}\neq 0\end{subarray}}^{B}\!\!\!\!\!\!\operatorname*{softmax}_{j}(G_{RbM}(\mathbf{h}_{j})_{i})\,\mathbf{h}_{j}\,,$ (4)
where $\hat{\lambda}(t)=\max(\lambda(t),1.0)$ is a coefficient. We use an annealing schedule to ensure that $\lambda(t)<\lambda(t+1)$. As $t\rightarrow\infty$, the change in the expert embeddings tends to zero, facilitating the convergence of all embeddings.
We compute this sum only for the vectors that are routed to the expert. Applying the softmax over the batch dimension improves performance.
@@
where $s\in\operatorname{\mathbb{R}}$ is a learnable output scaler, $\mathbf{att}_{i}\in\operatorname{\mathbb{R}}^{d}$ is a learnable input attention vector, and $\scalerel\mathbin{\scalerel*{\bm{\odot}}{\circ}}$ denotes element-wise multiplication. Note that the attention is applied only to the expert input and not to the router input. This routing approach is inspired by the technique presented by Zhang et al. (2021a); we extend it to perform top-k rather than top-1 routing and employ cosine similarity instead of Euclidean distance.
@@
$Loss_{VQ}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{E}G_{RbM}(\mathbf{h}_{i})_{j}\frac{\operatorname{sg}[\mathbf{Q}^{RbM}_{j}]\mathbf{h}_{i}}{\|\operatorname{sg}[\mathbf{Q}^{RbM}_{j}]\|\|\mathbf{h}_{i}\|}\,,$ (6)
@@
$Loss_{SS}=\frac{1}{E^{2}}\sum_{i=1}^{E}\sum_{j=1}^{E}\frac{\operatorname{sg}[\mathbf{Q}^{RbM}_{j}]\mathbf{Q}^{RbM}_{i}}{\|\operatorname{sg}[\mathbf{Q}^{RbM}_{j}]\|\|\mathbf{Q}^{RbM}_{i}\|}\,.$ (7)
@@
$Loss_{LB}=\frac{Var(\mathbf{Load})}{Mean(\mathbf{Load})^{2}},\quad\quad\textrm{for}\quad\mathbf{Load}=\sum_{i=1}^{B}G_{RbM}(\mathbf{h}_{i}).$ (8)
Here $\mathbf{Load}\in\operatorname{\mathbb{R}}^{E}$ is the vector of unnormalized expert utilization. While expert specialization does not require load balancing, we adopt this loss to redistribute input vectors that lie in almost equal proximity to multiple experts. This loss encourages representations to move closer to the least utilised expert embeddings.
@@ 4.2 Knowledge Distillation
$Loss_{KD}=\frac{\nu}{\left|\mathcal{V}^{L}\right|}\sum_{v\in\mathcal{V}^{L}}CE(\bf{\hat{y}}_{v},\bf{y}_{v})+\frac{1-\nu}{\left|\mathcal{V}\right|}\sum_{v\in\mathcal{V}}KL(\bf{\hat{y}}_{v},\bf{\hat{y}^{\prime}}_{v})\,.$ (9)
Here $\nu$ is a hyperparameter that controls how much the knowledge distillation training relies on the soft labels from the teacher GNN, $CE(\cdot,\cdot)$ denotes cross-entropy, and $KL(\cdot,\cdot)$ is a KL-divergence (Zhang et al., 2021b). While the supervised part of the loss uses only the labeled part of the graph, the KL-divergence is computed over the entire set of nodes.
@@
In order to additionally leverage the graph structure, we employ the knowledge-based sampling technique of KRD (Wu et al., 2023). KRD is based on the principle of aligning the representation of a target node with those of reliable nodes in its neighbourhood. To accomplish this, Wu et al. (2023) introduce a measure of reliability, which is based on identifying nodes whose teacher GNN representations do not change substantially when the features in the graph are subjected to perturbation. The reliability metric for node $j$ is:
$\rho_{j}=\frac{1}{\delta^{2}}\operatorname{\mathbb{E}}_{\mathbf{\tilde{X}}\sim N(\mathbf{X},\delta\mathbf{I})}\left\lVert\mathcal{H}(\mathbf{\hat{y}}^{\prime}_{j})-\mathcal{H}(\mathbf{\tilde{y}^{\prime}}_{j})\right\rVert^{2},$ (10)
where $N(\cdot,\delta\mathbf{I})$ is Gaussian noise with diagonal covariance matrix, with $\delta$ as the diagonal elements. $\mathcal{H}(\cdot)$ denotes the entropy, and $\mathbf{\tilde{y}}^{\prime}_{j}$ and $\mathbf{\mathbf{\hat{y}}}^{\prime}_{j}$ are the predicted class distributions for node $j$ with and without perturbation, respectively. This metric is based on the principle that more reliable teacher soft labels are more robust to feature perturbations.
During knowledge distillation, given a target node $v$, we sample a set of nodes from its neighbourhood. The sampling distribution is determined by the reliability metric:
$p(j|\rho_{j},\alpha)\propto 1-\left(\frac{\rho_{j}}{\rho_{\max}}\right)^{\alpha},$ (11)
where $\rho_{\max}=\operatorname*{arg\,max}_{i\in\mathcal{V}}\rho_{i}$, and $\alpha$ is a learnable power parameter. To learn $\alpha$, we follow the procedure of (Wu et al., 2023). During training of the overall model, we strive to match the class distribution predicted by the student $\mathbf{\hat{y}}_{v}$ with the teacher-provided soft labels of the sampled nodes. We therefore include a loss term:
$Loss_{KRD}=\frac{1-\nu}{\left|\mathcal{V}\right|}\sum_{v\in\mathcal{V}}\operatorname{\mathbb{E}}_{\begin{subarray}{c}u\in\mathcal{N}(v)\\ u\sim\bar{p}(u|\rho_{u},\alpha)\end{subarray}}KL(\mathbf{\hat{y}}_{v},\mathbf{\hat{y}}^{\prime}_{u}),$ (12)
where, for each node $v$, the expectation is with respect to a distribution $\bar{p}(u|\rho_{u},\alpha)$, which is proportional to $p(u|\rho_{u},\alpha)$ defined in equation 11, for the nodes $u$ in the neighborhood of $v$, and zero elsewhere. The hyperparameter $\nu$ is the same as in equation 9.
Our model consist of $L$ MoE layers (see Figure 3). The training objective is a combination of the cross-entropy loss, the distillation loss, the KRD loss, and embedding losses for every MoE layer of the model. With a model with $L$ MoE layers, we introduce weights $\alpha$, $\beta$, and $\gamma$ to determine the influences of the associated embedding losses: